# Streaming Generation of Co-Speech Gestures via Accelerated Rolling Diffusion

### v1

@@ Abstract
Generating co-speech gestures in real time requires both temporal coherence and efficient sampling. We introduce Accelerated Rolling Diffusion, a novel framework for streaming gesture generation that extends rolling diffusion models with structured progressive noise scheduling, enabling seamless long-sequence motion synthesis while preserving realism and diversity. We further propose Rolling Diffusion Ladder Acceleration (RDLA), a new approach that restructures the noise schedule into a stepwise ladder, allowing multiple frames to be denoised simultaneously. This significantly improves sampling efficiency while maintaining motion consistency, achieving up to a 2× speedup with high visual fidelity and temporal coherence. We evaluate our approach on ZEGGS and BEAT, strong benchmarks for real-world applicability. Our framework is universally applicable to any diffusion-based gesture generation model, transforming it into a streaming approach. Applied to three state-of-the-art methods, it consistently outperforms them, demonstrating its effectiveness as a generalizable and efficient solution for real-time, high-fidelity co-speech gesture synthesis.
@@ 1 Introduction
Co-speech gestures significantly enhance non-verbal communication by reinforcing spoken content, crucially contributing to realism in virtual avatars, video conferencing, gaming, and interactive embodied AI applications [24]. Real-time generation of these gestures, or streaming generation, is essential in scenarios like virtual assistants, gaming, and telepresence systems.
Recent approaches predominantly leverage data-driven deep learning methods, conditioned on audio, text, or style-specific attributes to produce realistic motion sequences, commonly represented in the standardized BioVision Hierarchy (BVH) format. While prior generative frameworks including GANs [11], VQ-VAEs [21], and flow-based models [36] have shown success, diffusion-based models have recently emerged as particularly effective due to their exceptional realism and diversity [34, 41, 8, 14].
Nevertheless, diffusion-based methods still face significant limitations for real-time scenarios. Typically, these methods generate fixed-length gesture sequences, stitching segments together via sliding windows, which can result in discontinuities and latency due to necessary post-processing [35]. Approaches enhancing continuity by conditioning current frames on previous ones introduce heavy computational overheads from expanding context windows, reducing their applicability in real-time interactive systems.
To address these challenges, we introduce Accelerated Rolling Diffusion, a novel framework integrating diffusion models with real-time capabilities for co-speech gesture generation. By employing a structured ladder-based noise scheduling strategy, our Rolling Diffusion Ladder Acceleration (RDLA) approach simultaneously denoises multiple frames, significantly improving sampling efficiency. This achieves generation speeds of up to 120 FPS without compromising visual fidelity or temporal coherence. Extensive experiments on the ZEGGS [9] and BEAT [20] benchmarks confirm our method’s effectiveness for realistic, diverse gesture generation in streaming contexts. In summary, our key contributions are:
@@
    We are the first, to our knowledge, to successfully adapt rolling diffusion framework to a practical application, specifically demonstrating their effectiveness in real-time co-speech gesture generation.
@@
    We propose a universal framework converting any diffusion-based gesture generation approach into a real-time streaming model without requiring post-processing.
@@
    We introduce RDLA, substantially improving inference speed with minimal impact on gesture quality.
@@
    We provide comprehensive evaluations on standard benchmarks and user studies, validating our approach’s efficiency, realism, and robustness in real-time applications.
2 Related Work
Long-sequence Motion Generation. Generating long gesture sequences is challenging due to variable context-driven lengths and memory constraints. Gestures are typically conditioned by text, music, or speech, each requiring tailored approaches. Text-based methods often generate segments individually, stitching them together using weakly supervised techniques with motion smoothness priors [22, 39], iterative refinement like TEACH [2], or diffusion-based methods (e.g., DoubleTake) for temporal consistency [29]. Precise temporal control is provided by multi-track timelines [25] and blended positional encodings [4]. Additionally, hierarchical models like MotionMamba [40] and MultiAct [17], along with language-driven frameworks such as MotionGPT [15], enable seamless, coherent generation across extended sequences.
Dancing Motions Generation. Long-sequence dance synthesis poses challenges due to duration and style diversity. Methods like EDGE [33] align overlapping clips for smooth transitions, while ”You Never Stop Dancing” [32] uses low-dimensional manifolds with RefineBank and TransitBank to maintain fluid, high-quality motion.
Co-speech Gesture Generation. Driven by the pursuit of natural and expressive human-computer interactions, co-speech gesture generation has rapidly advanced using deep generative techniques. Researchers have explored methods from GANs, VAEs, and VQ-VAEs to diffusion models enhanced with transformer attention. While GANs [11] often face instability and mode collapse, diffusion models deliver robustness, high fidelity, and diverse outputs. For instance, DiffuseStyleGesture [34] integrates cross-local and self-attention to synchronize varied gestures, and EMAGE [21] employs masked audio-conditioned modeling with VQ-VAEs for greater expressivity. TalkSHOW [37] separately generates facial, body, and hand motions for nuanced speech alignment, while DiffTED [14] uses TPS keypoints in a diffusion pipeline to improve coherence. Lastly, Audio to Photoreal Embodiment [23] combines diffusion with vector quantization to create realistic conversational avatars.
Long-sequence Co-Speech Gestures Generation Generating coherent long gesture sequences is challenging due to potential discontinuities and temporal inconsistencies. A common approach is to generate short chunks separately and then stitch overlapping clips together seamlessly [1, 21, 34, 41, 38]. For example, FreeTalker [35] employs the DoubleTake blending technique to ensure smooth transitions, while autoregressive methods condition each new gesture on preceding outputs to preserve temporal coherence. Additionally, frameworks like DiffSHEG [8] use outpainting strategies to extend sequences incrementally, and DiffTED [14] generates coherent TPS keypoint sequences that bridge gesture synthesis with realistic video rendering.
Diffusion Models have gained popularity for generative tasks, particularly in image and video synthesis. Early work introduced denoising score matching [31] and later evolved into denoising diffusion probabilistic models (DDPM) [12]. Their success in high-quality image synthesis, seen in models like Imagen [28] and Stable Diffusion [26], has been extended to video generation [13, 3, 18, 6]. Given their temporal consistency and ability to model complex distributions, diffusion models are also being explored for co-speech gesture generation.
Rolling Diffusion Models [27] extend traditional diffusion-based generation to sequential data, enabling autoregressive synthesis by iteratively generating and conditioning on previous outputs. This approach enhances temporal consistency and long-range dependencies. Recent works [7, 16, 10] continue to refine this concept making diffusion models more effective for sequential data generation by enhancing temporal consistency, enabling infinite-length generation, and improving temporal dynamics modeling.
@@ 3 Proposed Approach
3.1 Diffusion Models
Diffusion models consist of a forward (diffusion) process and reverse process. The forward process gradually adds Gaussian noise to a data sample $x^{0}$ over $T$ steps such that $x^{T}\sim\mathcal{N}(0,1)$ and each transition is defined as:
@@ 3.1 Diffusion Models
The reverse process learns to denoise $x^{t}$ using a neural network $f_{\theta}(x^{t},t)=\hat{x}$. The model is trained by minimizing the objective: $\mathcal{L}_{\theta}(x^{0},t):=\mathbb{E}_{t,\epsilon,\mathbf{x}}\left[a(t)\|x^{0}-\hat{x}\|^{2}\right]$, where $a(t)$ is a weighting function that can be specified to control the importance of different timesteps during training. This objective encourages $f_{\theta}(x^{t},t)$ to accurately estimate the initial signal of data sample.
3.2 Rolling Diffusion Models
Rolling Diffusion Models (RDMs) [27] introduce a modification to standard diffusion models by incorporating a progressive corruption process along the temporal axis, making them particularly well-suited for sequential data $\mathbf{X}=\{x^{0}_{l}\}_{l=0}^{L-1}$, where $x_{l}$ is $l$-th element of sequence. Unlike standard diffusion models that apply noise uniformly across all frames, Rolling Diffusion Models (RDMs) operate with a rolling window $\mathbf{x}_{j}=\{x_{j+n}\}_{n=0}^{N-1}$. In this setup, the noise level gradually increases from the first to the last frame in the window, enabling a seamless transition. Once the first frame is fully denoised, a new frame, sampled from Gaussian noise, is introduced, and the window shifts forward to continue the denoising process. This approach allows for continuous and unbounded generation, making RDMs particularly effective for producing arbitrarily long sequences.
@@ 3.2 Rolling Diffusion Models
$q(\mathbf{x}^{t}_{j}|\mathbf{x}_{j})=\prod_{n=0}^{N-1}\mathcal{N}(x^{t_{n}}_{j+n}|\alpha^{t_{n}}x_{j+n}^{0},(\sigma^{t_{n}})^{2}I)$ (3)
During training, the model $f_{\theta}(\mathbf{x}_{j}^{t},t)=\hat{\mathbf{x}}$ processes only the frames within the rolling window. The parametrized reverse process $p_{\theta}(\mathbf{x}^{t-1}_{j}|\mathbf{x}^{t}_{j})$ is defined as:
$p_{\theta}(\mathbf{x}^{t-1}_{j}|\mathbf{x}^{t}_{j})=\prod_{n=0}^{N-1}q(x_{j+n}^{t_{n}-1}|x_{j+n}^{t_{n}},\hat{x}_{j+n})$ (4)
The training objective:
$L_{\theta}(\mathbf{x}_{j},t):=\sum_{n=0}^{N-1}a(t_{n})\|x^{0}_{j+n}-\hat{x}_{j+n}\|^{2}$ (5)
where $x_{n}^{t}$ represents a single frame with the upper index $t$ denoting the noise level and the lower index $n$ representing the index in the sequence.
@@
3.3 Method
In our work, we adapt rolling diffusion models for co-speech gesture generation, introducing a novel framework that transforms any diffusion-based architecture into a streaming model. Our approach enables seamless and continuous gesture generation of arbitrary length by modifying the model architecture and integrating a structured noise scheduling mechanism, which, combined with the rolling denoising process, ensures smooth temporal transitions and prevents abrupt motion discontinuities. In our implementation, we discretize time using $t\in[0,T]$ instead of the continuous range $t\in[0,1]$, where $T=1000$ represents the total number of noise steps. Since the generation window size $N$ is typically much smaller than $T$, the noise level difference between adjacent frames is greater than 1. We define this difference as a step $s$, calculated as $s=\frac{T}{N}$. To ensure uniform noise distribution across frames, we select the generation window as a divisor of $T$, ensuring consistent step sizes and preventing uneven noise application. This structured noise scheduling allows for a more controlled and stable generation process, improving the overall quality of generated sequences. As a condition, the model receives audio features as input $U=\{u_{l}\}_{l=0}^{L-1}$. Audio processing depends on the implementation of the baseline model, but the most popular approach is to use a pre-trained model such as WavLM. Also, depending on the baseline’s architecture, the model can receive a speech style or speaker ID as an additional condition.
A key modification in our framework is replacing the baseline model’s single time-conditioning value per sequence with a time sequence matching the generation window’s length. This enables the model to handle varying noise levels across frames within the window. However, to minimize architectural changes, especially when the model incorporates time embeddings through concatenation instead of summation, we can employ a single time value. This value represents the noise level of the first frame. Given that each subsequent frame within a fixed time window experiences progressively higher noise, their corresponding noise levels are effectively determined by this initial time value.
3.4 Training description
@@
During training, we sample the initial noise level $t_{0}$ for the first frame from a uniform distribution $t_{0}\sim\text{Uniform}(\{1,\dots,s\})$ and select the starting index $j$ for the rolling window uniformly $j\sim\text{Uniform}(\{1,\dots,L-N\})$. We then determine the noise levels for subsequent frames according to $t_{n}=t_{0}+s\cdot n$. These noise levels are applied using Eq. 3. As a result, the noise level for each frame $n$ in the sequence falls within the range $t_{n}\in[s\cdot n,\;s\cdot(n+1))$. To improve the performance of the model, we include several clean frames at the beginning of the sequence as an additional context, represented by $\mathbf{x}_{j}^{cont}=(x_{j-n^{cont}}^{1},\dots,x_{j-1}^{1})$ with a length $n^{cont}$. We discovered that applying a minimal level of noise $t=1$ to these context frames is essential. This process acts as a form of regularization and improves stability. (See Supplemental Material for more details). Therefore, at each time step, the input to the model consists of a concatenated sequence $[\mathbf{x}^{cont}_{j},\mathbf{x}^{t}_{j}]$. Here, $\mathbf{x}^{t}_{j}$ denotes a rolling window whose first frame corresponds to the $t$-th noise level, and the associated audio features for this sequence are given by $u_{j}=(u_{j-n^{cont}},\dots,u_{j+N-1})$.
@@
Unlike prior work, our model is trained exclusively for the rolling phase, omitting an initial descent phase. Furthermore, we use $a(t)=1$ in the training objective for all $t$, instead of using the signal-to-noise ratio (SNR). This modification reduces complexity while maintaining stable training dynamics. The training procedure is summarized in Algorithm 1.
@@
4:     $\mathbf{x}_{j}\leftarrow(x^{0}_{j},\dots,x^{0}_{j+N-1})$
@@
7:     Compute $[\hat{\mathbf{x}}^{\mathrm{cont}}_{j},\,\hat{\mathbf{x}}_{j}]\leftarrow f_{\theta}\Bigl{(}[\mathbf{x}^{\mathrm{cont}}_{j},\,\mathbf{x}^{t_{0}}_{j}],\,t_{0},\,\mathbf{u}_{j}\Bigr{)}$
@@
3.5 Sampling process description
@@
1:audio $\{u_{j}\}_{j=0}^{L-1}$, idle $x_{\mathrm{idle}}$, resulted prediction $y$
2:for $n=-N,\dots,-1$ do
3:     $t_{n}=s(N+n+1)$
4:     $u_{n}=0$
5:     Sample $x_{n}^{t_{n}}\sim q(x_{n}^{t_{n}}|x_{\mathrm{idle}})$
6:end for
7:$\mathbf{u}_{-N}\leftarrow(u_{-N-n^{cont}},\dots,u_{-1})$
8:$\mathbf{x}_{-N}^{s}\leftarrow(x_{-N}^{s},\dots,x_{-1}^{T})$
9:$\mathbf{x}_{-N}^{\mathrm{cont}}\leftarrow({x}_{\mathrm{idle}},\dots,x_{\mathrm{idle}})$
10:$j=-N$
11:repeat
12:     for $t_{0}=s,s-1,...,1$ do
13:         $[\hat{\mathbf{x}}^{\mathrm{cont}}_{j},\,\hat{\mathbf{x}}_{j}]\leftarrow f_{\theta}\Bigl{(}[\mathbf{x}^{\mathrm{cont}}_{j},\,\mathbf{x}^{t_{0}}_{j}],\,t_{0},\,\mathbf{u}_{j}\Bigr{)}$
14:         Sample $\mathbf{x}^{t_{0}-1}_{j}\sim p_{\theta}(\mathbf{x}^{t_{0}-1}_{l}|\mathbf{x}^{t_{0}}_{j})$ (see Eq. 4)
15:     end for
16:     $y_{j}=x_{j}^{0};j=j+1$
17:     Sample $x_{j-1}^{1}\sim q(x_{j-1}^{1}|x_{j-1}^{0});x^{T}_{N}\sim\mathcal{N}(0,I)$
18:     $\mathbf{x}_{j}^{cont}\leftarrow(x^{1}_{j-n^{\mathrm{cont}}},\dots,x_{j-1}^{1})$
19:     $\mathbf{x}_{j}^{s}\leftarrow(x_{j}^{s},\dots,x_{j+N}^{T})$
20:     $\mathbf{u}_{j}\leftarrow(u_{j-n^{cont}},\dots,u_{j+N})$
21:until Completed

### v3

@@ Abstract
Generating co-speech gestures in real time requires both temporal coherence and efficient sampling. We introduce a novel framework for streaming gesture generation that extends Rolling Diffusion models with structured progressive noise scheduling, enabling seamless long-sequence motion synthesis while preserving realism and diversity. Our framework is universally compatible with existing diffusion-based gesture generation model, transforming them into streaming methods capable of continuous generation without requiring post-processing. We evaluate our framework on ZEGGS and BEAT, strong benchmarks for real-world applicability. Applied to state-of-the-art baselines on both datasets, it consistently outperforms them, demonstrating its effectiveness as a generalizable and efficient solution for real-time co-speech gesture synthesis. We further propose Rolling Diffusion Ladder Acceleration (RDLA), a new approach that employs a ladder-based noise scheduling strategy to simultaneously denoise multiple frames. This significantly improves sampling efficiency while maintaining motion consistency, achieving up to a 4× speedup with high visual fidelity and temporal coherence in our experiments. Comprehensive user studies further validate our framework’s ability to generate realistic, diverse gestures closely synchronized with the audio input.
Codehttps://github.com/andrewbo29/co-speech-gestures-rolling-diffusion
@@ 1 Introduction
Co-speech gestures enhance non-verbal communication by reinforcing spoken content, and are essential for realism in virtual avatars, video conferencing, gaming, and embodied AI (nyatsanga2023comprehensive). Real-time, or streaming, generation is particularly important for interactive settings such as virtual assistants, gaming, and telepresence systems. Recent work largely relies on data-driven deep learning to produce realistic motion, with diffusion-based models standing out for their strong realism and diversity. Approaches such as DiffuseStyleGesture (yang2023diffusestylegesture), which uses cross-local and self-attention for gesture–speech synchronization, and Taming Diffusion (zhu2023tamingdiffusionmodelsaudiodriven), which applies audio-conditioned diffusion with transformer guidance, both rely on seed-pose conditioning to ensure smooth transitions and temporal coherence.
Despite recent progress, diffusion-based methods still struggle in real-time settings. To manage long sequences, many models generate fixed-length chunks or incremental extensions. PersonaGestor (zhang2024speechdrivenpersonalizedgesturesynthetics) uses fuzzy feature inference yet still stitches chunked outputs, while DiffSHEG (chen2024diffsheg) depends on incremental outpainting. Such non-streaming strategies often cause visual discontinuities and add latency due to post-processing. Methods that condition on previous frames improve continuity but incur substantial computational overhead, limiting real-time use (see Supplementary Material).
A promising alternative is the Rolling Diffusion framework (ruhe2024rollingdiffusionmodels), which turns diffusion models into an autoregressive process and improves temporal consistency for long sequences. However, applying it to real-time co-speech gesture generation remains difficult, primarily because the autoregressive loop is computationally expensive.
To address these challenges, we introduce a novel Rolling Diffusion based framework integrating diffusion models with real-time capabilities for co-speech gesture generation. By employing a structured ladder-based noise scheduling strategy, our Rolling Diffusion Ladder Acceleration (RDLA) approach simultaneously denoises multiple frames, significantly improving sampling efficiency. This achieves generation speeds of up to 200 FPS without compromising visual fidelity or temporal coherence. Extensive experiments on the ZEGGS (ghorbani2023zeroeggs) and BEAT (liu2022beatlargescalesemanticemotional) benchmarks confirm our method’s effectiveness for realistic, diverse gesture generation in streaming contexts. In summary, our key contributions are:
@@
    We are the first, to our knowledge, to successfully adapt a rolling diffusion framework to a practical application, specifically demonstrating its effectiveness in real-time co-speech gesture generation.
@@
    We propose a universal framework for converting any diffusion-based gesture generation approach into a real-time streaming model without requiring post-processing.
@@
    We provide comprehensive evaluations on standard benchmarks and user studies, showing that our method consistently outperforms baselines across the benchmarks, demonstrating its effectiveness as a generalizable solution for real-time, high-fidelity co-speech gesture synthesis.
@@
    We introduce RDLA, substantially increasing inference speed with minimal impact on gesture quality.
@@ 2 Proposed Approach
Rolling Diffusion Models
Diffusion models (song2020generativemodelingestimatinggradients; ho2020denoising) consist of a forward (diffusion) process and reverse process. The forward process gradually adds Gaussian noise to a data sample $x^{0}$ over $T$ steps such that $x^{T}\sim\mathcal{N}(0,1)$ and each transition is defined as:
@@
The reverse process learns to denoise $x^{t}$ using a neural network $f_{\theta}(x^{t},t)=\hat{x}$. The model is trained by minimizing the objective: $\mathcal{L}(\theta):=\mathbb{E}_{t,x^{0}}\left[a(t)\|x^{0}-\hat{x}\|^{2}\right]$, where $a(t)$ is a weighting function that can be specified to control the importance of different timesteps during training. This objective encourages $f_{\theta}(x^{t},t)$ to accurately estimate the initial signal of data sample.
Rolling Diffusion Models (RDMs) (ruhe2024rollingdiffusionmodels) introduce a modification to standard diffusion models by incorporating a progressive corruption process along the temporal axis, making them particularly well-suited for sequential data $\mathbf{X}=\{x^{0}_{k}\}_{k=0}^{L-1}$, where $x_{l}$ is $k$-th element of sequence. Unlike standard diffusion models that apply noise uniformly across all frames, RDMs operate with a rolling window of size $N$. In this setup, the noise level gradually increases from the first to the last frame in the window, enabling a seamless transition. In our implementation, we discretize time which indicates the noise level using $t\in\{0,\ldots,T\}$ instead of the continuous range $t\in[0,1]$, where $T=1000$ represents the total number of noise steps. Since the generation window size $N$ is typically much smaller than $T$, the noise level difference between adjacent frames is greater than 1. We define this difference as a step $s$, calculated as $s=\frac{T}{N}$. To ensure uniform noise distribution across frames, we select the generation window as a divisor of $T$, ensuring consistent step sizes and preventing uneven noise application. This structured noise scheduling allows for a more controlled and stable generation process, improving the overall quality of generated sequences. Then the rolling window takes the form $\mathbf{x}^{t_{0}}_{j}=\{x_{j+n}^{t_{n}}\}_{n=0}^{N-1}$, where $t_{0}\in\{1,\ldots,s\}$ is a noise level of the first frame in the window, $t_{n}=t_{0}+ns$ is a noise level of $n^{th}$ frame.  [^1] Since noise levels of all frames within a rolling window are deterministic functions of $t_{0}$ we may use this index to define the noise levels of the entire rolling window $\mathbf{x}_{j}^{t_{0}}$.
@@
$q(\mathbf{x}^{t_{0}}_{j}|\mathbf{x}_{j})=\prod_{n=0}^{N-1}\mathcal{N}(x^{t_{n}}_{j+n}|\alpha^{t_{n}}x_{j+n}^{0},(1-\alpha^{t_{n}})^{2}I)$ (3)
During training, the model $f_{\theta}(\mathbf{x}_{j}^{t_{0}},t)=\hat{\mathbf{x}}$ processes only the frames within the rolling window. The parametrized reverse process $p_{\theta}(\mathbf{x}^{t-1}_{j}|\mathbf{x}^{t}_{j})$ is defined as:
$p_{\theta}(\mathbf{x}^{t_{0}-1}_{j}|\mathbf{x}^{t_{0}}_{j})=\prod_{n=0}^{N-1}q(x_{j+n}^{t_{n}-1}|x_{j+n}^{t_{n}},\hat{x}_{j+n})$ (4)
Figure 1: Visualization of the rolling denoising process with parameters $T=5$, $N=5$, $n^{cont}=1$, $s=1$
Once the first frame is fully denoised, i.e. $t_{0}$ reaches 0, a new frame, sampled from Gaussian noise, is introduced, and the window shifts forward ($j=j+1$, $t_{0}=s$) to continue the denoising process. This approach allows for continuous and unbounded generation, making RDMs particularly effective for producing arbitrarily long sequences.
The training objective is
$L(\theta):=\mathbb{E}_{j}\mathbb{E}_{\mathbf{x}^{0}_{j}}\sum_{t^{0}=1}^{s}\mathbb{E}_{\mathbf{x}_{j}^{t_{0}}}\sum_{n=0}^{N-1}a(t_{n})\|x^{0}_{j+n}-\hat{x}_{j+n}\|^{2},$ (5)
where $x_{k}^{t}$ represents a single frame with the upper index $t$ denoting the noise level and the lower index $k$ representing the index in the sequence and $a(t_{n})$ is a weighting function (e.g. signal-to-noise ratio).
@@
Method
In our work, we adapt rolling diffusion models for co-speech gesture generation, introducing a novel framework that transforms any diffusion-based architecture into a streaming model. Our approach enables seamless and continuous gesture generation of arbitrary length by modifying the model architecture and integrating a structured noise scheduling mechanism, which, combined with the rolling denoising process, ensures smooth temporal transitions and prevents abrupt motion discontinuities. As illustrated in figure 1, the model generates a new clean frame in each $s$-step and shifts the generation window forward to include the new frame at the end.
As a condition, the model receives audio features as input $U=\{u_{k}\}_{k=0}^{L-1}$. Audio processing depends on the implementation of the baseline model, but the most popular approach is to use a pre-trained model such as WavLM (chen2022wavlm). Also, depending on the baseline’s architecture, the model can receive a speech style or speaker ID as an additional condition.
The architecture of the baseline model is mostly preserved, and the main change affects the integration of time embedding. All models incorporate time embeddings into each frame, either by summing or concatenating them with the frame embeddings. In the original models, a single time embedding is replicated across all frames. In contrast, our approach assigns a unique time embedding to each frame within the rolling window.
Training Description
@@
During training, we sample the initial noise level $t_{0}$ for the first frame from a uniform distribution $t_{0}\sim\text{Uniform}(\{1,\dots,s\})$ and select the starting index $j$ for the rolling window uniformly $j\sim\text{Uniform}(\{1,\dots,L-N\})$. We then determine the noise levels for subsequent frames according to $t_{n}=t_{0}+ns$. These noise levels are applied using eq. 3. As a result, the noise level for each frame $n$ in the sequence falls within the range $t_{n}\in[s\cdot n,\;s\cdot(n+1))$. To improve the performance of the model, we include several clean frames at the beginning of the sequence as an additional context, represented by $\mathbf{x}_{j}^{cont}=(x_{j-n^{cont}}^{1},\dots,x_{j-1}^{1})$ with a length $n^{cont}$. We discovered that applying a minimal level of noise $t=1$ to these context frames is essential. This process acts as a form of regularization and improves stability (see the Supplementary Material for more details). Therefore, at each time step, the input to the model consists of a concatenated sequence $\mathbf{x}^{cont}_{j},\mathbf{x}^{t}_{j}]$. Here, $\mathbf{x}^{t}_{j}$ denotes a rolling window whose first frame corresponds to the $t$-th noise level, and the associated audio features for this sequence are given by $u_{j}=(u_{j-n^{cont}},\dots,u_{j+N-1})$.
@@
Unlike prior work, our model is trained exclusively for the rolling phase, omitting an initial descent phase. Furthermore, we use $a(t_{n})=1$ in the training objective for all $t$, instead of using the signal-to-noise ratio (SNR). The training procedure is summarized in algorithm 1.
@@
4:  $\mathbf{x}^{0}_{j}\leftarrow(x^{0}_{j},\dots,x^{0}_{j+N-1})$
@@
7:  Compute $\hat{\mathbf{x}}_{j}\leftarrow f_{\theta}\Bigl([\mathbf{x}^{\mathrm{cont}}_{j},\,\mathbf{x}^{t_{0}}_{j}],\,t_{0},\,\mathbf{u}_{j}\Bigr)$
@@
Sampling Process Description
@@
1:audio $\{u_{j}\}_{j=0}^{L-1}$, idle $x_{\mathrm{idle}}$
2:Resulted prediction $y$
3:for $n=-N,\dots,-1$ do
4:  $t_{n}=s(N+n+1)$
5:  $u_{n}=0$
6:  Sample $x_{n}^{t_{n}}\sim q(x_{n}^{t_{n}}|x_{\mathrm{idle}})$
7:end for
[…]
11:j=Nj=-N
12:repeat
13:  for $t_{0}=s,s-1,...,1$ do
14:   $\hat{\mathbf{x}}_{j}\leftarrow f_{\theta}\Bigl([\mathbf{x}^{\mathrm{cont}}_{j},\,\mathbf{x}^{t_{0}}_{j}],\,t_{0},\,\mathbf{u}_{j}\Bigr)$
15:   Sample $\mathbf{x}^{t_{0}-1}_{j}\sim p_{\theta}(\mathbf{x}^{t_{0}-1}_{l}|\mathbf{x}^{t_{0}}_{j})$ (see eq. 4)
16:  end for
17:  $y_{j}=\hat{x}_{j}^{0};j=j+1$
18:  Sample $x_{j-1}^{1}\sim q(x_{j-1}^{1}|x_{j-1}^{0});x^{T}_{j+N}\sim\mathcal{N}(0,I)$
19:  $\mathbf{x}_{j}^{cont}\leftarrow(x^{1}_{j-n^{\mathrm{cont}}},\dots,x_{j-1}^{1})$
20:  $\mathbf{x}_{j}^{s}\leftarrow(x_{j}^{s},\dots,x_{j+N}^{T})$
21:  $\mathbf{u}_{j}\leftarrow(u_{j-n^{cont}},\dots,u_{j+N})$
22:until Completed